An executive drove from home at an average speed of 30 mph to an airport where a helicopter was waiting. The executive boarded the helicopter and flew to the corporate offices at an average speed of 80 mph. The entire distance was 135 mi. The entire trip took 2 h. Find the distance from the airport to the corporate offices.
Answer by Hawksfan(61)

t1 = t = the time it took to go from home to the airport where the helicopter was
t2 = 2-t = the time starting at the airport and ending at the corporate offices
r1 = 30
r2 = 80
d = 135
Using the d= rt formula and putting in what you know:
We need to find t:
d = r1 * t1 + r2 * t2
135 = 30 *t + 80 (2-t)
135 = 30t + 160 -80t
-25 =-50t
t = 0.5 h
Now we aren't finished because the question asked for the distance between the
airport and the corporate offices. We have a figure of time.
we need to find d2 = r2 * t2. We know t2 = 2-t. t2 = 2-0.5 = 1.5 h

d2 = r2 * t2 = 80 * 1.5 = 120mi
The distance is 120 mi